# Littlewood, Paley and Almost-Orthogonality: a theory well ahead of its time

### v1

@@ 10. Reprise: Almost-everywhere Fourier convergence in $\mathbb{R}^{n}$ for $n\geq 2$
In this final section we return to the pointwise convergence problems which originally motivated Littlewood and Paley in [LP1], but now in their higher-dimensional form.
There are various settings in which we can consider almost-everywhere Fourier convergence – and the corresponding maximal function – in dimensions $n\geq 2$. One of the most attractive of these is in euclidean space $\mathbb{R}^{n}$ where we consider the spherical partial inverse Fourier integrals (10) for $f$ belonging to various $L^{p}$ spaces, or perhaps subspaces consisting of functions with minimal smoothness. For $f\in L^{2}$ there is no problem in posing the questions since the Fourier transform is an isometric isomorphism of $L^{2}$.
For $p>2$ there are issues with even making sense of the questions. Indeed, unless $S_{R}$ maps the Schwartz space $\mathcal{S}(\mathbb{R}^{n})$ into $L^{p^{\prime}}(\mathbb{R}^{n})$, it is difficult to see how to define $S_{R}f$ even as a tempered distribution for all $f$ belonging to $L^{p}$. The integral kernel of $S_{R}$ can be calculated explicitly, and it is in $L^{p^{\prime}}$ if and only if $p^{\prime}>2n/(n+1)$. So the question for $L^{p}$ for $p>2$ only makes sense in the range p<2n/(n1)p<2n/(n-1), and even then we have to take care. These considerations also underly the occurrence of the index 2n/(n1)2n/(n-1) in (11).
For $p<2$, almost-everywhere convergence of $S_{2^{k}}f$ to $f$ for all $f\in L^{p}$ would imply via Nikišin–Stein factorisation theory (see [GCRdeF]) that the maximal operator $f\mapsto\sup_{k\in\mathbb{Z}}|S_{2^{k}}f|$ would be of weak-type $(p,p)$, and thus $S_{1}$ would be bounded on some $L^{q}$ for $q<2$, which would contradict Fefferman’s disc multiplier theorem [Feffermanball].
@@
Part (i) is in [CRdeFV]. The corresponding maximal function estimate for $p=2$ is folklore and can be proved by the same arguments as Theorem 2.3 part (i). Part (ii) is in [CarberySoria] and for $p=2$ is essentially a statement of Rademacher–Menšov type. As a corollary of part (ii) we have almost-everywhere convergence for $f$ in any Sobolev space $L^{p}_{s}$ of strictly positive order $s$ when $2\leq p<2n/(n-1)$. It is not currently known whether one may recover the result of part (i) when one replaces the lacunary sequence $\{2^{k}\}_{k\in\mathbb{Z}}$ by a second-order lacunary sequence $\{2^{k}(1+2^{-j})\}_{k\in\mathbb{Z},j\in\mathbb{N}}$ (this corresponds to the maximal Hilbert transform in the one-dimensional setting). However, almost-everywhere convergence through a second-order lacunary sequence does hold if $\int_{\mathbb{R}^{n}}|\widehat{f}(\xi)|^{2}(1+\log^{+}\log^{+}|\xi|)^{2}{\rm d}\xi<\infty$, see [CarbGMT].

### v2

@@ 10. Reprise: Almost-everywhere Fourier convergence in $\mathbb{R}^{n}$ for $n\geq 2$
In this final section we return to the pointwise convergence problems which originally motivated Littlewood and Paley in [50], but now in their higher-dimensional form.
There are various settings in which we can consider almost-everywhere Fourier convergence – and the corresponding maximal function – in dimensions $n\geq 2$. One of the most attractive of these is in euclidean space $\mathbb{R}^{n}$ where we consider the spherical partial inverse Fourier integrals (10) for $f$ belonging to various $L^{p}$ spaces, or perhaps to subspaces consisting of functions with minimal smoothness. For $f\in L^{2}$ there is no problem in posing the questions since the Fourier transform is an isometric isomorphism of $L^{2}$.
For $p>2$ there are issues with even making sense of the questions. Indeed, unless $S_{R}$ maps the Schwartz space $\mathcal{S}(\mathbb{R}^{n})$ into $L^{p^{\prime}}(\mathbb{R}^{n})$, it is difficult to see how to define $S_{R}f$ even as a tempered distribution for all $f$ belonging to $L^{p}$. The integral kernel of $S_{R}$ can be calculated explicitly, and it is in $L^{p^{\prime}}$ if and only if $p^{\prime}>2n/(n+1)$. So the question for $L^{p}$ for $p>2$ only makes sense, even in principle, in the range p<2n/(n1)p<2n/(n-1), and even then we have to take care. These considerations also underly the occurrence of the index 2n/(n1)2n/(n-1) in (11).
For $p<2$, almost-everywhere convergence of $S_{2^{k}}f$ to $f$ for all $f\in L^{p}$ would imply via Maurey–Nikišin–Stein factorisation theory (see [37]) that the maximal operator $f\mapsto\sup_{k\in\mathbb{Z}}|S_{2^{k}}f|$ would be of weak-type $(p,p)$, and thus $S_{1}$ would be bounded on some $L^{q}$ for $q<2$, which would contradict Fefferman’s disc multiplier theorem [35].
@@
Part (i) is in [13]. The corresponding maximal function estimate for $p=2$ is folklore and can be proved by the same arguments as Theorem 2.3 part (i). Part (ii) is in [14] and for $p=2$ is essentially a statement of Rademacher–Menšov type. It would be interesting to lower the index $2$ on the logarithmic term to $1$. As a corollary of part (ii) we have almost-everywhere convergence for $f$ in any Sobolev space $L^{p}_{s}$ of strictly positive order $s$ when $2\leq p<2n/(n-1)$. It is not currently known whether one may recover the result of part (i) when one replaces the lacunary sequence $\{2^{k}\}_{k\in\mathbb{Z}}$ by a second-order lacunary sequence $\{2^{k}(1+2^{-j})\}_{k\in\mathbb{Z},j\in\mathbb{N}}$ (this corresponds to the maximal Hilbert transform in the one-dimensional setting). However, almost-everywhere convergence through a second-order lacunary sequence does hold if $\int_{\mathbb{R}^{n}}|\widehat{f}(\xi)|^{2}(1+\log^{+}\log^{+}|\xi|)^{2}{\rm d}\xi<\infty$, see [12].